# An Effective Version of the $p$-Curvature Conjecture for Order One Differential Equations

### v1

@@ 4.2.1 Choice of the Number of Functions to Approximate
For any $M>\frac{1}{2}(C_{0}\Delta^{3}\delta^{3}-1)$ we have $X(M)<1$. Moreover, for $M>C_{0}\Delta^{3}\delta^{3}$ we have $X(M)<2^{-(2M+1)}(C_{0}\Delta\delta)^{-2}$.
Доказательство.
@@
The choices of $M$ and $N$ are dependent on each other in the following way: the function $X(M)$ is decreasing with $M$, while simultaneously $Y(M,N)$ increases. Thus, the first value of $N$, such that $X(M)^{N}Y(M,N)Z(M,N)<1$ varies in a non-obvious way with $M$. To obtain tight bounds one should minimize $\sigma=(2M+1)N+2M$ under the constraint $X(M)^{N}Y(M,N)Z(M,N)<1$. We carry out no such computations, but are contempt with $M=\lceil C_{0}\Delta^{3}\delta^{3}\rceil$. Naïve estimates and computations suggest that picking $M\approx\lceil 0.806\cdot C_{0}\Delta^{3}\delta^{3}\rceil$ yields slightly more optimal values for $\sigma$.
@@ 4.2.2 Choice of the Type in the Approximation
The goal of this section is to prove that for the explicit values of $M$ and $N$ given in Theorem 1.2 the right-hand side of (14) is bounded from above by $1$. We assume $M=\lceil C_{0}\Delta^{3}\delta^{3}\rceil$ to be fixed. In Proposition 4.7, we assume $N$ to depend at least linearly on $B$ and $M$. Let us set $N\coloneqq ABM$ for some constant $A\geq 4$ to be determined later. With these choices of $M$ and $N$ the right-hand side of (14) is a function in $\Delta$. However, to simplify computations we will treat it as a function in $M=\lceil C_{0}\Delta^{3}\delta^{3}\rceil$.
@@
For $M=\lceil C_{0}\Delta^{3}\delta^{3}\rceil$ and $N=ABM$ with $A\geq 4$, we have
@@
For $A>6.076$ and all $B\geq 1$ and $\Delta\geq 1$ the expression $T(M)$ is negative and thus
@@
Доказательство.
We estimate crudely $\frac{2e^{2}\Delta^{5}\delta^{2}}{N\pi^{2}}<M$. Moreover, by Lemma 4.11, $X(M)<(C_{0}\Delta\delta)^{-2}\cdot 2^{-(2M+1)}$. We have $\log(C_{0}^{2}\Delta^{2}\delta^{2})=-\frac{4}{3}\log(C_{0})-\frac{2}{3}\log(C_{0}\Delta^{3}\delta^{3})$, and $\log(C_{0}\Delta^{3}\delta^{3})\leq-\log(M-1)\leq-\log(M)-\log(\frac{2}{3})$ by factoring $M$ in the $\log$ and using that $M\geq 3$, hence $\log(C_{0}^{2}\Delta^{2}\delta^{2})\leq-\frac{2}{3}\log(\frac{2C_{0}^{2}}{3})-\frac{2}{3}\log(M)$. We further use $\log(2M+1)>\log(M)$ and bound $\frac{M+1}{6N}<\frac{1}{18}$. The rest of the proof of the first part is obtained from the definitions of $X(M)$ and $Y(M,N)$ by straightforward computations.
@@ 5 Algorithm and Complexity
The computation of the normal form, as described above, of $u(x)\in\mathbb{Q}(x)$ of height $h$ and degree n1n-1 in the numerator and $n$ in the denominator requires computing the greatest common divisor of $n+1$ numbers and the greatest common divisor of two polynomials. The computation of the greatest common divisor of two polynomials in $\mathbb{Q}[x]$ of degree at most $n$ and height $h$ can be made by performing $\tilde{O}(nh)$ bit operations [vzGG13, Corollary 11.9, p.325]. With an adaptation of the same algorithm we can compute the gcd of two integers of height at most $h$ in $\tilde{O}(h)$ bit operations. Hence applying it $n$ times allows to compute the greatest common divisor of $n+1$ integers of height at most $h$ in $\tilde{O}(nh)$ bit operations, thus computing the normal form of $u(x)$ can be made by performing $\tilde{O}(nh)$ bit operations.
@@
6:while $p\leq\Delta$ do
7:  if $\Delta\bmod p=0$ then $\delta\leftarrow\delta\cdot p^{1/(p-1)}$
8:  else if $\texttt{pCurvature}(a(x),b(x),p)\neq 0$ then return Transcendental;
@@
10:Compute $B$, $M\leftarrow\lceil 2.826\Delta^{3}t^{3}\rceil$, $N\leftarrow\lceil 6.076BM\rceil$;
@@
As explained in Remark 3.1, starting from $u(x)\in\overline{\mathbb{Q}}(x)$, we can check if $\tilde{R}(w)\coloneqq R(w)/\Delta$ is in $\mathbb{Q}[w]$ after step 4. If this is not the case we can immediately conclude that the output has to be Transcendental with Proposition 2.6. Otherwise, we take $\Delta$ to be the discriminant of the polynomial in $\mathbb{Z}[w]$ associated to $\tilde{R}(w)$.
@@ 5.1 Complexity of Rothstein-Trager Resultants
Let $u(x)=a(x)/b(x)\in\mathbb{Q}(x)$ be in normal form with $\deg a(x)<n\coloneqq\deg b(x)$, with the denominator $b(x)$ being squarefree, height bounded by $h$ and multiplicative height at most $H=2^{h}$.
@@
Доказательство.
The coefficient of $x^{n-1}$ in $a(x)-w\cdot b^{\prime}(x)$ cannot vanish hence $\deg_{x}(a(x)-w\cdot b^{\prime}(x))=n-1$. Corollary 11.21 from [vzGG13] concludes. ∎
@@
From the form of this Sylvester matrix, we see that the degree of $R(w)$ is at most $n$, and its coefficient of degree $n$ is $\operatorname{res}(b(x),b^{\prime}(x))$, up to a sign. The assumption that $b(x)$ is squarefree ensures $\deg(R(w))=n$. Concerning the height of $R(w)$ we prove the following estimate.
@@
The main ingredient for our estimates is the following lemma known as Hadamard’s inequality [vzGG13, Thm. 16.6, p. 477].
@@ 5.2 Complexity Estimates for Finding Rational Roots of the Rothstein-Trager Resultant
In Section 3.1 we noted algebraicity of the solutions of (2) can be decided by Algorithm 1. We briefly investigate the complexity of this algorithm. For efficiently finding all the rational roots of a polynomial we have the following complexity [Bos+17, Prop. 21.22].
@@
Let $R(w)\in\mathbb{Z}[w]$ be a squarefree, primitive polynomial of degree $n$ and height $h$. The computation of all rational roots of $R(w)$ can be performed, in $\tilde{O}(n^{2}h)$ bit operations.
@@
Von zur Gathen and Gerhard only give a probabilistic algorithm performing the task in the same expected complexity [vzGG13, Thm. 15.21]. The algorithm described in [Bos+17] uses a deterministic (potentially costly) algorithm to find roots of the polynomial modulo a deterministically defined small prime number, and then lifts them to characteristic zero. It improves a result by Loos [Loo83], who proposed a deterministic algorithm for finding rational roots in $\tilde{O}(n^{3}h)$ bit operations.
@@
Given $u(x)\coloneqq\frac{a(x)}{b(x)}\in\mathbb{Q}(x)$ of degree $n$ and height $h$, deciding if all solutions of $y^{\prime}(x)=u(x)y(x)$ are algebraic can be done by performing $\tilde{O}(n^{2}h)$ bit operations.
@@ 5.3 Computation of the Bound on the Primes
Given $u(x)$, and having computed the Rothstein-Trager resultant $R(w)\in\mathbb{Z}[w]$, Theorem 1.2 gives an explicit bound $\sigma\in\mathbb{N}$ on the number of $p$-curvatures to check, depending on $\Delta$, the leading coefficient, and an upper bound $B$ on the modulus of all roots of $R(w)$.
An arbitrarily precise estimate of $B$ is possible without factoring completely the polynomial $R(w)$, for example by following ideas from an unpublished report of Schцnhage in which an algorithm to find the complex roots of a polynomial $R(w)\in\mathbb{C}[w]$ with arbitrary precision is presented [Sch82, Thm. 15.1].
@@
Before computing a $p$-curvature, we need to ensure that $p\not\hskip 2.5pt\mid\Delta$, hence we can find the prime factors of $\Delta$ step by step, without factoring $\Delta$. After checking all $p$-curvatures for $p\leq\Delta$, we can compute $\delta$, then $M$, NN and $\sigma$.
@@ 5.4 Checking the Vanishing of $p$-Curvatures
To compute $p$-curvatures we rely on an algorithm due to Bostan and Schost [BS09] which is tailored specifically for first-order differential equations. By ‘‘computing’’ a $p$-curvature, we mean that we compute enough terms of its Taylor expansion in order to uniquely reconstruct it knowing a bound on the degrees of its numerator and denominator. Indeed, a rational function with numerator’s and denominator’s degrees at most $n$ is completely determined by $n$ and the first $2n$ coefficients of its Taylor series expansion.
@@
The computation of the expansion of $\bar{u}(x)$ by performing a Newton iteration can be done in $\tilde{O}(n\log(p))$ bit operations [vzGG13, Thm. 9.4, p. 260]. If we write $\bar{u}(x)=\bar{u}_{0}+\bar{u}_{1}x+\dots+\bar{u}_{i}\in\mathbb{F}_{p}$, then the Taylor expansion of $\bar{v}(x)$ is $\bar{u}_{p-1}+\bar{u}_{2p-1}x+\bar{u}_{3p-1}x^{2}+\dots$ The extraction of these specific coefficients can be done with an algorithm due to Fiduccia [Fid85], performing $\tilde{O}(n\log(p)(n+\log(p)))$ bit operations. This algorithm uses the following result proved by Fiduccia: if we write $f(x)\coloneqq x^{n-\operatorname{val}_{x}(b(x))}\bar{b}(\frac{1}{x})\in\mathbb{F}_{p}[x]$ for the characteristic polynomial of the sequence $(\bar{u}_{i})$, then for any $k\in\mathbb{N}_{\geq n}$, $\bar{u}_{k}=\varphi_{0,k}\bar{u}_{0}+\dots+\varphi_{n-1,k}\bar{u}_{n-1}$ where $\varphi_{i,k}\in\mathbb{F}_{p}$ are defined by $x^{k}=\varphi_{0,k}+\varphi_{1,k}x+\dots+\varphi_{n-1,k}x^{n-1}\bmod f(x)$. Knowing the first $2n$ terms $\bar{u}_{0},\dots,\bar{u}_{2n-1}$ of the Taylor expansion of $\bar{u}(x)$, the computation of the p1p-1-th term $\bar{u}_{p-1}$ amounts to compute a power of the image $\xi$ of $x$ in the ring $\mathbb{F}_{p}[x]/(f(x))$ and this can be made by binary powering in $\tilde{O}(n\log(p)^{2})$ bit operations. Now that we have computed $\xi^{p-1}$, it only takes one product in $\mathbb{F}_{p}[x]/(f(x))$ to compute $\xi^{p}$, and one more for each $\xi^{(i+1)p-1}$ knowing $\xi^{ip-1}$. In total computing the coefficients of $v(x)$ up to order $2n$ takes $\tilde{O}(n\log(p)(n+\log(p)))$ with this method.
@@
An efficient algorithm to compute the characteristic polynomials of the $p$-curvatures for any order differential equations for all primes $p\leq S$ simultaneously in $\tilde{O}(S(\log(H)+n)n^{\omega+1})$ bit operations is given by Pagиs in [Pag21], where $\omega<2.3728596$ is an exponent for matrix multiplication. This is easily adapted to check the nilpotence of the $p$-curvatures in the same complexity. Checking nullity in such a complexity could prove crucial in the development of efficient algorithms for a potential effective version of the Grothendieck $p$-curvature conjecture for (classes of) higher order equations.

### v2

@@ 4.2.1 Choice of the Number of Functions to Approximate
For any $M>\frac{1}{2}(C_{0}r_{n}^{3}\delta^{3}-1)$ we have $X(M)<1$. Moreover, for $M>C_{0}r_{n}^{3}\delta^{3}$ we have $X(M)<2^{-(2M+1)}(C_{0}r_{n}\delta)^{-2}$.
Proof.
@@
The choices of $M$ and $N$ are dependent on each other in the following way: the function $X(M)$ is decreasing with $M$, while simultaneously $Y(M,N)$ increases. Thus, the first value of $N$, such that $X(M)^{N}Y(M,N)Z(M,N)<1$ varies in a non-obvious way with $M$. To obtain tight bounds one should minimize $\sigma=(2M+1)N+2M$ under the constraint $X(M)^{N}Y(M,N)Z(M,N)<1$. We carry out no such computations, but are contempt with $M=\lceil C_{0}r_{n}^{3}\delta^{3}\rceil$. Naïve estimates and computations suggest that picking $M\approx\lceil 0.806\cdot C_{0}r_{n}^{3}\delta^{3}\rceil$ yields slightly more optimal values for $\sigma$.
@@ 4.2.2 Choice of the Type in the Approximation
The goal of this section is to prove that for the explicit values of $M$ and $N$ given in Theorem 1.2 the right-hand side of (14) is bounded from above by $1$. We assume $M=\lceil C_{0}r_{n}^{3}\delta^{3}\rceil$ to be fixed. In Proposition 4.7, we assume $N$ to depend at least linearly on $B$ and $M$. Let us set $N\coloneqq ABM$ for some constant $A\geq 4$ to be determined later. With these choices of $M$ and $N$ the right-hand side of (14) is a function in $r_{n}$ and $\Delta_{R}$. However, to simplify computations we will treat it as a function in $M=\lceil C_{0}r_{n}^{3}\delta(\Delta_{R})^{3}\rceil$.
@@
For $M=\lceil C_{0}r_{n}^{3}\delta^{3}\rceil$ and $N=ABM$ with $A\geq 4$, we have
@@
For $A>6.076$ and all $B\geq 1$ and $r_{n}\geq 1$ the expression $T(M)$ is negative and thus
@@
Proof.
We estimate crudely $\frac{2e^{2}r_{n}^{5}\delta^{2}}{N\pi^{2}}<M$. Moreover, by Lemma 4.11, $X(M)<(C_{0}r_{n}\delta)^{-2}\cdot 2^{-(2M+1)}$. We have $\log(C_{0}^{2}r_{n}^{2}\delta^{2})=-\frac{4}{3}\log(C_{0})-\frac{2}{3}\log(C_{0}r_{n}^{3}\delta^{3})$, and $\log(C_{0}r_{n}^{3}\delta^{3})\leq-\log(M-1)\leq-\log(M)-\log(\frac{2}{3})$ by factoring $M$ in the $\log$ and using that $M\geq 3$, hence $\log(C_{0}^{2}r_{n}^{2}\delta^{2})\leq-\frac{2}{3}\log(\frac{2C_{0}^{2}}{3})-\frac{2}{3}\log(M)$. We further use $\log(2M+1)>\log(M)$ and bound $\frac{M+1}{6N}<\frac{1}{18}$. The rest of the proof of the first part is obtained from the definitions of $X(M)$ and $Y(M,N)$ by straightforward computations.
@@ 5 Algorithm and Complexity
The computation of the normal form, as described above, of $u(x)\in\mathbb{Q}(x)$ of height $h$ and degree n1n-1 in the numerator and $n$ in the denominator requires computing the greatest common divisor of $n+1$ numbers and the greatest common divisor of two polynomials. The computation of the greatest common divisor of two polynomials in $\mathbb{Q}[x]$ of degree at most $n$ and height $h$ can be made by performing $\tilde{O}(nh)$ bit operations [13, Corollary 11.9, p.325]. With an adaptation of the same algorithm we can compute the gcd of two integers of height at most $h$ in $\tilde{O}(h)$ bit operations. Hence applying it $n$ times allows to compute the greatest common divisor of $n+1$ integers of height at most $h$ in $\tilde{O}(nh)$ bit operations, thus computing the normal form of $u(x)$ can be made by performing $\tilde{O}(nh)$ bit operations.
Additionally, we will assume in the following that the Rothstein-Trager resultant $R(w)=\operatorname{res}_{x}(b(x),a(x)-w\cdot b^{\prime}(x))$ is square-free to lighten the exposition. If it is not, we replace it with its square-free part, which can be readily computed as $R(w)/\gcd(R(w),R^{\prime}(w))$ in negligible bit operation cost. The polynomial $R(w)$ and its square-free part have the same roots, hence the same bound $B$ on their modulus, and the same prime factors of the leading coefficient. However, the height of $R(w)$ might be altered when taking the square-free part. We prove an estimate in Proposition 5.7 that ensures the order of the height remains the same.
Last, to avoid costly integer factoring of the potentially large resultant $\Delta_{R}$ to compute $\delta$, we prove an estimate in Proposition 5.15.
@@
[…]
7:while $p\leq\Delta_{b}$ do
8:  if $\Delta_{b}\bmod p\neq 0$ and $\texttt{pCurvature}(a(x),b(x),p)\neq 0$ then return Transcendental;
@@
10:Compute $B$, $M\leftarrow\lceil 2.826\Delta_{b}^{3}\delta^{3}\rceil$, $N\leftarrow\lceil 6.076BM\rceil$;
@@
As explained in Remark 3.1, starting from $u(x)\in\overline{\mathbb{Q}}(x)$, we can check if $\tilde{R}(w)\coloneqq R(w)/\Delta_{b}$ is in $\mathbb{Q}[w]$ after step 4. If this is not the case we can immediately conclude that the output has to be Transcendental with Proposition 2.6. Otherwise, we can continue from step 5 after replacement of $R(w)$ by the polynomial in $\mathbb{Z}[w]$ associated to $\tilde{R}(w)$ obtained by multiplication by the denominator of the coefficients.
@@ 5.1 Complexity of Rothstein-Trager Resultants
Let $u(x)=a(x)/b(x)\in\mathbb{Q}(x)$ be in normal form with $\deg a(x)<n\coloneqq\deg b(x)$, with the denominator $b(x)$ being square-free, height bounded by $h$ and multiplicative height at most $H=2^{h}$.
@@
Proof.
The coefficient of $x^{n-1}$ in $a(x)-w\cdot b^{\prime}(x)$ cannot vanish hence $\deg_{x}(a(x)-w\cdot b^{\prime}(x))=n-1$. Corollary 11.21 from [13] concludes. ∎
@@
From the form of this Sylvester matrix, we see that the degree of $R(w)$ is at most $n$, and its coefficient of degree $n$ is $\operatorname{res}(b(x),b^{\prime}(x))$, up to a sign. The assumption that $b(x)$ is square-free ensures $\deg(R(w))=n$. Concerning the height of $R(w)$ we prove the following estimate.
@@
The main ingredient for our estimates is the following lemma known as Hadamard’s inequality [13, Thm. 16.6, p. 477].
@@
Taking the square-free part of the Rothstein-Trager resultant can alter its height, let us bound the height of any of its factors.
Proposition 5.7.
Any factor of $R(w)=\operatorname{res}_{x}(b(x),a(x)-w\cdot b^{\prime}(x))$ has height at most
$\frac{2^{2n+1}}{\pi n}\cdot H^{2n-1}\cdot(n+1)^{2n}.$
Proof.
Let $H_{R}$ be the height of $R(w)\coloneqq r_{n}w^{n}+\dots+r_{0}$. Let $\|R(w)\|_{2}=\left(\sum_{k=0}^{n}|r_{k}|^{2}\right)^{1/2}$. Assume $P(w)\in\mathbb{Z}[w]$ is a factor of $R(w)$, Theorem 2 in [31] ensures that for the height $H_{P}$ of $P(w)$ we have
$H_{P}\leq\max\left(\binom{\ell}{k},\ 0\leq k\leq\ell\leq n\right)\|R(w)\|_{2}.$
We have naïvely $\|R(w)\|_{2}\leq(n+1)^{1/2}H_{R}$. The maximum for fixed $\ell$ is the central binomial coefficient $\binom{\ell}{\lfloor\ell/2\rfloor}$, which is an increasing sequence of $\ell$, thus $H_{P}\leq\binom{n}{\lfloor n/2\rfloor}(n+1)^{1/2}H_{R}$. Combining this with Proposition 5.3 we obtain $H_{P}\leq(n+1)^{2n}\binom{n}{\lfloor n/2\rfloor}^{2}H^{2n-1}$. The following upper bound on the central binomial coefficient yields the announced result
$\binom{n}{\lfloor n/2\rfloor}\leq\frac{2^{n+1/2}}{\sqrt{\pi n}}.$
If $n$ is even this is a classical result, let us prove that it holds for any odd $n\in\mathbb{N}_{\geq 3}$. We have
$\binom{n}{\frac{n-1}{2}}=\binom{n-1}{\frac{n-1}{2}}+\binom{n-1}{\frac{n-3}{2}}\leq\frac{2^{n-1/2}}{\sqrt{\pi(n-1)}}\cdot\left(1+\frac{n-1}{n+1}\right).$
The result follows by dividing by the desired bound and using that $x\mapsto\frac{1}{2}\left(\frac{x}{x-1}\right)^{1/2}\left(1+\frac{x-1}{x+1}\right)$ attains its minimum $\frac{3\sqrt{3}}{4\sqrt{2}}<1$ on $2,+\infty)$ at $x=3$. ∎
@@ 5.2 Complexity Estimates for Finding Rational Roots of the Rothstein-Trager Resultant
In Section 3.1 we noted algebraicity of the solutions of (2) can be decided by Algorithm 1. We briefly investigate the complexity of this algorithm. For efficiently finding all the rational roots of a polynomial we have the following complexity [7, Prop. 21.22].
@@
Let $R(w)\in\mathbb{Z}[w]$ be a square-free, primitive polynomial of degree $n$ and height $h$. The computation of all rational roots of $R(w)$ can be performed, in $\tilde{O}(n^{2}h)$ bit operations.
@@
Von zur Gathen and Gerhard only give a probabilistic algorithm performing the task in the same expected complexity [13, Thm. 15.21]. The algorithm described in [7] uses a deterministic (potentially costly) algorithm to find roots of the polynomial modulo a deterministically defined small prime number, and then lifts them to characteristic zero. It improves a result by Loos [28], who proposed a deterministic algorithm for finding rational roots in $\tilde{O}(n^{3}h)$ bit operations.
@@
Given $u(x)\coloneqq\frac{a(x)}{b(x)}\in\mathbb{Q}(x)$ of degree $n$ and height $h$, deciding if all solutions of $y^{\prime}(x)=u(x)y(x)$ are algebraic can be done by performing $\tilde{O}(n^{3}h)$ bit operations.
@@ 5.3 Computation of the Bound on the Primes
Given $u(x)$, and having computed the Rothstein-Trager resultant $R(w)\in\mathbb{Z}[w]$, Theorem 1.2 gives an explicit bound $\sigma\in\mathbb{N}$ on the number of $p$-curvatures to check, depending on $r_{n}$, the leading coefficient, on $\Delta_{R}\coloneq\operatorname{res}_{w}(R(w),R^{\prime}(w))$, and on an upper bound $B$ on the modulus of all roots of $R(w)$.
An arbitrarily precise estimate of $B$ is possible without factoring completely the polynomial $R(w)$, for example by following ideas from an unpublished report of Schцnhage in which an algorithm to find the complex roots of a polynomial $R(w)\in\mathbb{C}[w]$ with arbitrary precision is presented [45, Thm. 15.1].
@@
Before computing a $p$-curvature, we check if $p|r_{n}$. After checking that all $p$-curvatures vanish for $p\leq\Delta_{b}$, we can compute $\delta$, then $M$, NN and $\sigma$.
@@ 5.4 Checking the Vanishing of $p$-Curvatures
To compute $p$-curvatures we rely on an algorithm due to Bostan and Schost [9] which is tailored specifically for first-order differential equations. By ‘‘computing’’ a $p$-curvature, we mean that we compute enough terms of its Taylor expansion in order to uniquely reconstruct it knowing a bound on the degrees of its numerator and denominator. Indeed, a rational function with numerator’s and denominator’s degrees at most $n$ is completely determined by $n$ and the first $2n$ coefficients of its Taylor series expansion.
@@
The computation of the expansion of $\bar{u}(x)$ by performing a Newton iteration can be done in $\tilde{O}(n\log(p))$ bit operations [13, Thm. 9.4, p. 260]. If we write $\bar{u}(x)=\bar{u}_{0}+\bar{u}_{1}x+\dots+\bar{u}_{i}\in\mathbb{F}_{p}$, then the Taylor expansion of $\bar{v}(x)$ is $\bar{u}_{p-1}+\bar{u}_{2p-1}x+\bar{u}_{3p-1}x^{2}+\dots$ The extraction of these specific coefficients can be done with an algorithm due to Fiduccia [14], performing $\tilde{O}(n\log(p)(n+\log(p)))$ bit operations. This algorithm uses the following result proved by Fiduccia: if we write $f(x)\coloneqq x^{n-\operatorname{val}_{x}(b(x))}\bar{b}(\frac{1}{x})\in\mathbb{F}_{p}[x]$ for the characteristic polynomial of the sequence $(\bar{u}_{i})$, then for any $k\in\mathbb{N}_{\geq n}$, $\bar{u}_{k}=\varphi_{0,k}\bar{u}_{0}+\dots+\varphi_{n-1,k}\bar{u}_{n-1}$ where $\varphi_{i,k}\in\mathbb{F}_{p}$ are defined by $x^{k}=\varphi_{0,k}+\varphi_{1,k}x+\dots+\varphi_{n-1,k}x^{n-1}\bmod f(x)$. Knowing the first $2n$ terms $\bar{u}_{0},\dots,\bar{u}_{2n-1}$ of the Taylor expansion of $\bar{u}(x)$, the computation of the p1p-1-th term $\bar{u}_{p-1}$ amounts to compute a power of the image $\xi$ of $x$ in the ring $\mathbb{F}_{p}[x]/(f(x))$ and this can be made by binary powering in $\tilde{O}(n\log(p)^{2})$ bit operations. Now that we have computed $\xi^{p-1}$, it only takes one product in $\mathbb{F}_{p}[x]/(f(x))$ to compute $\xi^{p}$, and one more for each $\xi^{(i+1)p-1}$ knowing $\xi^{ip-1}$. In total computing the coefficients of $v(x)$ up to order $2n$ takes $\tilde{O}(n\log(p)(n+\log(p)))$ with this method.
@@
An efficient algorithm to compute the characteristic polynomials of the $p$-curvatures for any order differential equations for all primes $p\leq S$ simultaneously in $\tilde{O}(S(\log(H)+n)n^{\omega+1})$ bit operations is given by Pagиs in [37], where $\omega<2.3728596$ is an exponent for matrix multiplication. This is easily adapted to check the nilpotence of the $p$-curvatures in the same complexity. Checking nullity in such a complexity could prove crucial in the development of efficient algorithms for a potential effective version of the Grothendieck $p$-curvature conjecture for (classes of) higher order equations.